Dain’s Diamond Bit Drilling purchased the following assets this year. (2018)

Purchase Original
Asset Date Basis
Drill bits (5-year) Feb-22 $ 88,500
Drill bits (5-year) Sep-12 101,250
Commercial building Apr-12 251,000

Assume its taxable income for the year was $60,000 for purposes of computing the §179 expense (assume no bonus depreciation). 2018

c. If the February drill bits’ original basis was $2,384,000, what is the maximum amount of §179 expense Dain may deduct for the year?

§179 expense
Answer #1

Maximum amount of $179 expense is 2384000-88500=2295500*10months/12=$1912917

The total amount of loses is compensate on the proportion of months vested by Dian holding.